Status Report

ISSP Vehicle Health 6 Dec 2002

By SpaceRef Editor
January 14, 2003
Filed under , ,

Source: ISS Monthly Program Review 6 Dec 2002

Click to enlarge

Note: Worsened or Red status issues shown-new status in bold, underlined font


  • 2B and 4B BGA showing high currents sporadically

    • 2B and 4B rotated as required for power
    • Dual angle mode and XPOP flown when possible
    • Low b X-POP being pursued to limit BGA rotations, when possible.

  • 3 of 4 Beta Gimbal Assembly (BGA) latching mechanisms locked
    on starboard 4 Bar assembly/Strength analysis shows 3 of 4 acceptable for life.

  • Contamination on Node1 nadir CBM may impact future ops. Cleaning planned for
    Stage EVA


  • Early external active thermal control system operating
    within specs

    • Starboard radiator has one loop plumbed incorrectly/Heat rejection
    • capability impacted – still meets current heat rejection needs

  • ITCS operating nominally
    • ITCS fine filter R&R plan established.
    • Comprehensive ITCS fluid forward plan still in work


  • CanadArm2 operating nominally. RWS operational.


  • CMG #1 failed/All remaining CMGs have experienced occasional loss of comm
    • Momentum well within current requirements, no issue until 12A
    • Spare/Replacement scheduled for ULF-1

  • GPS system deactivated due to NaNdata sending GNC MDM to diagnostic mode.
    • Workaround in place.
    • Russian attitude data is prime.


  • Suspect PGTs replaced during 9A


  • No major anomalies


  • S-band high/low data rate operating nominally/Ku band operating nominally
  • MCOR operational
  • SM Regul System – 2 of 3 strings operational
  • SMKurs set #2 has intermittent failures – no impact


  • VOA (air sample analyzer) – sporadically operational
  • TOCA (water sample analyzer) – operational
  • TEPC (real-time radiation monitor) – Determined failed. Returned on 11A.
    Long-term plan in work.


  • TVIS – operating restrictions in place
    • Chassis R&R’d on 9A

  • CEVIS – Automatic mode drop-outs. Evaluating using manual mode for EVA


  • FGB EPS working nominally
    • 5 of 6 batteries on-line

  • SM EPS working nominally
    • All 8 batteries on-line

  • P6 power channels 2B and 4B operating nominally
  • 2B and 4B rotated as required for power
  • RPCMs
    • RPCMs LAD22B-A, LAFWD-1B-A, LAFWD-1B-C, and LAD62B-A
    • have bit flips on SRAM and cannot be refreshed
    • No short term impact
    • Spares available, if required
    • Overall R&R plan being implemented to return suspect RPCMs
    • to the ground


  • Lab ECLS systems operating as required
    • CDRA operational in dual bed mode. (Valves replaced on 11A)
    • FOD likely suspect for OCT problems. Parts will be inspected post-landing. Future plans TBD.
    • MCA operational.

  • Node smoke detector #2 operational, but off due to dirty indication –
  • Still awaiting on-orbit troubleshooting
  • Cleaning procedure in work.
  • Airlock CCAA water separator R&R’d but 9A stage restrictions in place.
  • No constraint for 11A EVA operations.
  • SMVozdukh operating on 2 of 3 CO2 removal beds after minor repairs pre-

  • CO2 removal capability nominal
  • SM Elektron Fluid Unit repaired


  • All MDMs operational

  • SpaceRef staff editor.